OpenStudy (anonymous):

a) Find all solutions of the equation. b) Find the solutions in the interval [0, 2pi),

OpenStudy (anonymous):

\[2\sin3\theta+1=0\]

OpenStudy (agreene):

a is impossible. in normal terms, because there are infinite solutions.

OpenStudy (agreene):

2 sin (3 theta) +1 = 0 sin (3 theta) = -1/2 sin^-1 both sides 3 theta = -pi/6 \[\theta = -\frac{\pi}{18}\] All solutions can be given in general terms as: \[\theta=\frac{\pi}{18}(12n-1)\space\space\space n\in\mathbb{Z}\] And \[\theta=\frac{\pi}{18}(12n+7)\space\space\space n\in\mathbb{Z}\]